Hello, Please help with these. I have to write an equation for the inverse of:
1. y= x+4 over 2
2. y=-1/2x + 2
3. Y=2x-3
Thank you so much!

Answer by lwsshak3(11628)

1. y= x+4 over 2
2. y=-1/2x + 2
3. Y=2x-3
**
To find the inverse(y^-1) of a one-to-one function, interchange x and y and solve for y:
1. y= x+4 over 2
x=(y+4)/2
2x=y+4
y^-1=2x-4
..
2. y=-1/2x + 2
x=-y/2+2
2x=-y+4
y^-1=-2x+4
..
3. Y=2x-3
x=2y-3
2y=x+3
y^-1=(x+3)/2